# Improving the Language Understanding Capabilities of Large Language Models Using Reinforcement Learning

### v4

@@ Appendix E Impact of Fine-Tuning on Language Modeling Ability
We experiment with SFT and PPO to improve NLU capabilities of LLMs and observe improved performance using PPO. However, it is crucial to ensure that fine-tuning methods do not significantly degrade the models’ general language generation abilities. To assess this, we directly evaluate the PPL (jel, 1977; Chelba and Jelinek, 2000) of LLAMA2-7B-chat-hf models fine-tuned on the SST-2 dataset using the WikiText-2 test set (Merity et al., 2016), which follows a natural human-written text distribution. We compare these fine-tuned models against the original, non-fine-tuned baseline model, with the expectation that the PPL of the fine-tuned models should closely match the baseline. Our results reveal that the original LLAMA2-7B-chat-hf achieves a perplexity of 6.939. The PPO-fine-tuned model closely maintains this baseline performance with a perplexity of 6.966, indicating minimal impact on its general language modeling capabilities. In contrast, the SFT-fine-tuned model displays a notably higher perplexity of 7.384, suggesting a significant reduction in generation capabilities due to convergence toward task-specific training distributions. We conjecture that PPO’s clipping mechanism effectively constrains policy updates, preventing large deviations from the reference model and thereby preserving the original language modeling capabilities of LLMs. These findings underscore PPO’s effectiveness in maintaining the general language abilities of LLMs during fine-tuning.
@@
Appendix F Comparison of RL Algorithms: PPO vs. GRPO
Algorithm SST-2 MRPC (F1) RTE CoLA QNLI Avg. Per-Step Runtime (s)
SFT 73.8 85.8 80.4 50.7 93.6 76.9 4.124
PPO 96.4 89.4 84.3 59.9 93.2 84.6 4.299
GRPO 96.7 91.2 88.5 55.2 93.1 84.9 5.155
Table 9: Zero-shot GLUE performance (SST-2, MRPC F1, RTE, CoLA, QNLI) and per-step runtime. Best score in each column is bolded.
Our objective is to improve the natural language understanding capabilities of the base (policy) model through RL fine-tuning. In this context, we compare two approaches: PPO and Group Relative Policy Optimization (GRPO) (Shao et al., 2024). PPO is highly effective but introduces additional computational overhead. This overhead stems from the need for repeated sampling and from updating a separate critic model to compute value functions. In contrast, GRPO was designed to mitigate these costs by bypassing the critic model entirely. Instead of maintaining a separate value network, GRPO samples multiple trajectories per prompt and computes each trajectory’s advantage by comparing its reward to the batch’s average (and standard deviation). This method not only simplifies the architecture but also reduces memory usage.
As detailed in Table 9, both PPO and GRPO deliver notable performance improvements over SFT. Notably, PPO only incurs about a 4% increase in per-step runtime compared to SFT. However, GRPO’s need to generate multiple responses per sample results in a higher runtime, despite its memory efficiency benefits. Overall, our analysis highlights the trade-offs between these RL algorithms: PPO offers efficient runtime with the cost of additional overhead from the critic model, while GRPO reduces memory usage at the expense of increased sampling time.
Appendix G Reward Function Design and Evaluation

### v5

@@ 3.7 Impact of Fine-Tuning on Language Modeling Ability
We experiment with SFT and PPO to improve NLU capabilities of LLMs and observe improved performance using PPO. However, it is crucial to ensure that fine-tuning methods do not significantly degrade the models’ general language generation abilities. To assess this, we directly evaluate the PPL (jel, 1977; Chelba and Jelinek, 2000) of LLAMA2-7B-chat-hf models fine-tuned on the SST-2 dataset using the WikiText-2 test set (Merity et al., 2016), which follows a natural human-written text distribution. We compare these fine-tuned models against the original, non-fine-tuned baseline model, with the expectation that the PPL of the fine-tuned models should closely match the baseline. Our results reveal that the original LLAMA2-7B-chat-hf achieves a perplexity of 6.939. The PPO-fine-tuned model closely maintains this baseline performance with a perplexity of 6.966, indicating minimal impact on its general language modeling capabilities.
In contrast, the SFT-fine-tuned model displays a higher perplexity of 7.384, suggesting a significant reduction in generation capabilities due to convergence toward task-specific training distributions. We conjecture that PPO’s clipping mechanism and the KL-divergence minimization term in the objective function (part of the TRL library), effectively constrains policy updates, preventing large deviations from the reference model and thereby preserving the original language modeling capabilities of LLMs. These findings underscore PPO’s effectiveness in maintaining the general language abilities of LLMs during fine-tuning.
@@
3.8 Integrating a Reward Model
@@
4 Conclusion
Limitations
This work takes an initial step toward framing NLU as a reinforcement learning problem for LLMs under 14B parameters. While our long-term goal is to reduce reliance on curated datasets by leveraging richer, task-specific reward models, we currently adopt a simple binary reward signal based on exact label matching. This design enables a controlled evaluation of PPO as an effective adaptation strategy, showing consistent gains over prompting and supervised fine-tuning. Although we present a preliminary exploration of model-driven reward functions in Section 3.8, further research is needed to develop robust and generalizable reward signals that can support more complex or weakly supervised tasks without requiring extensive manual annotation. Overall, our findings suggest that casting nuanced tasks as reinforcement learning problems, through the design of appropriate environments and reward functions, offers a scalable and flexible alternative to standard fine-tuning, particularly when the model is already well-initialized through pretraining.
Appendix A Related Works
Appendix B Preliminaries on Application of PPO for Fine-tuning LLMs
Proximal policy optimization (PPO) (Schulman et al., 2017b) is an online reinforcement learning algorithm. In this section, we describe the process to fine-tune an LLM using PPO. During training, at each timestep $t$, the LLM (policy) generates a token prediction $a_{t}$ (action) based on the state $s_{t}$, which consists of the sequence of generated tokens up to timestep t1t-1. The final generated output is evaluated in the context of the downstream task, where the environment provides feedback in the form of rewards. The model updates its parameters based on these rewards to improve its ability to generate accurate predictions over time.
Actor Model
$\begin{split}J(\theta)&=\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta^{\prime}}}\Biggl[\min\Bigl(\,\frac{p_{\theta}(a_{t}\mid s_{t})}{p_{\theta^{\prime}}(a_{t}\mid s_{t})}\,\hat{A}_{t},\\ &\qquad\quad\mathrm{clip}\!\bigl(\tfrac{p_{\theta}(a_{t}\mid s_{t})}{p_{\theta^{\prime}}(a_{t}\mid s_{t})},\,1-\epsilon,\,1+\epsilon\bigr)\hat{A}_{t}\Bigr)\Biggr]\end{split}$ (2)
Here, $p_{\theta}(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ in state $s_{t}$ under the current policy, while $p_{\theta^{\prime}}(a_{t}|s_{t})$ represents this probability under the old policy. In PPO, the training data—specifically, the state-action pairs $(s_{t},a_{t})$—are sampled using the old policy $\pi_{\theta^{\prime}}$ (the LLM before it is updated), rather than the new policy currently being optimized. Thus, the ratio $\frac{p_{\theta}(a_{t}|s_{t})}{p_{\theta^{\prime}}(a_{t}|s_{t})}$ accounts for how much the new policy has changed relative to the old policy and adjusts the likelihood of an action accordingly. This ratio is multiplied by $\hat{A}_{t}$, the Generalized Advantage Estimation (GAE) (Schulman et al., 2018), which measures how much better or worse an action $a_{t}$ is compared to the expected outcome under the current policy.
$\hat{A}_{t}=R_{t}+\gamma V_{t+1}-V_{t}+\gamma\lambda\hat{A}_{t+1},$
Critic Model
$\delta_{t}=(R_{t}+\gamma V_{t+1}-V_{t})^{2},$ (3)
We implement PPO using TRL library Face that includes a KL divergence term in the PPO objective 2.
Appendix C Hyperparameter Settings
For few-shot evaluations, we conducted a sweep over 1–5 shots on the validation set and reported results using the best-performing configuration for each experiment. We ensured that the total input length, including few-shot demonstrations, remained within the model’s context window (e.g., in MultiRC, where input lengths can be long). For our experiments, we utilized the TRL library  (von Werra et al., 2020) on a single Nvidia A100 GPU.
Tasks LLAMA2-7B PPO-ST LLAMA2-13B PPO-ST GPT-4o
Sentiment Analysis
Financial PhraseBank (96.2, 98.1) (96.9, 98.5) (96.6, 98.4)
Labelled Financial News (66.1, 74.6) (69.0, 76.6) (63.2, 72.2)
Mental Health (66.6, 67.7) (66.0, 67.1) (59.3, 60.5)
Emotion (77.4, 78.6) (75.8, 77.0) (77.0, 78.2)
Natural Language Inference
Babi-nli (64.3, 71.5) (65.1, 73.0) (58.8, 67.6)
SIGA-nli (39.0, 53.9) (40.6, 53.7) (28.5, 42.2)
Appendix D Reward Curve for PPO Fine-Tuning
Appendix E Confidence Intervals on Validation Set of GLUE Benchmark
Our GLUE results are obtained from the official online evaluation servers, which impose a submission limit and keep test-set labels hidden, so applying a bootstrapping procedure directly is not feasible. We compute 95% confidence intervals on the validation sets of the GLUE benchmark (available offline) using the following procedure (used in the paper). For each example in a given dataset, we generate 100 independent predictions. We then compute the evaluation metric over the entire dataset for each set of predictions, yielding a distribution of metric values. The 2.5th and 97.5th percentiles of this distribution define the bounds of our 95% confidence interval. The resulting intervals are reported in the table 9.
Dataset PPO (Eval Metric) PPO (95% CI) SFT (Eval Metric) SFT (95% CI)
CoLA 64.54 (59.24, 70.08) 60.10 (54.79, 65.13)
MRPC 88.41 (86.92, 91.00) 85.28 (84.24, 87.19)
STS-B 88.57 (87.07, 89.98) 85.68 (83.85, 86.72)
QQP 83.34 (82.83, 83.73) 79.76 (79.22, 80.17)
QNLI 93.17 (92.50, 93.88) 93.67 (92.90, 94.35)
MNLI-matched 88.63 (88.05, 89.30) 87.02 (86.32, 87.64)
MNLI-unmatched 88.95 (88.28, 89.53) 87.27 (86.54, 87.89)
RTE 84.90 (80.26, 89.64) 81.95 (77.56, 86.95)
SST-2 96.10 (94.56, 97.54) 96.44 (94.95, 97.37)
WNLI 74.65 (63.30, 85.87) 66.20 (49.30, 71.83)
Table 9: Validation set results on GLUE benchmark. Metrics: CoLA (Matthews corr.), MRPC & QQP (F1), STS-B (Spearman corr.), all others (accuracy). 95% confidence intervals are shown in parentheses.
Appendix F Results on SuperGLUE Benchmark
Models BoolQ CB COPA MultiRC ReCoRD RTE
BERT-large 77.4 75.7/83.6 70.6 70.0/24.0 72.0/71.3 71.6
BERT-large++ 79.0 84.7/90.4 73.8 70.0/24.1 72.0/71.3 79.0
LLAMA2-7B-chat-hf
Zero-shot prompting 75.8 26.4/43.6 57.0 51.9/20.3 27.0/26.2 59.2
Few-shot prompting 80.2 49.8/66.0 73.4 46.6/15.4 36.3/35.3 72.9
PPO-ST 85.9 74.7/88.0 88.6 82.5/50.0 70.6/69.9 84.3
Models WiC WSC AXb AXg Average
BERT-large 69.5 64.3 23.0 97.8/51.7 69.0
BERT-large++ 69.5 64.3 38.0 99.4/51.4 71.5
LLAMA2-7B-chat-hf
Zero-shot prompting 54.4 52.1 9.1 64.0/55.1 49.5
Few-shot prompting 54.4 62.3 9.1 64.0/55.1 54.9
PPO-ST 72.1 78.1 52.7 91.0/79.8 78.3
We fine-tuned the LLAMA2-7B-chat-hf model using PPO on the SuperGLUE dataset and compared its performance against several baselines, including BERT-large, BERT-large++, and zero-shot and few-shot prompting of LLAMA2-7B-chat-hf. The term “BERT++” refers to a BERT model fine-tuned using the supplementary training on intermediate labeled-data tasks (STILTs) approach (Phang et al., 2018), where the model is first fine-tuned on related transfer tasks before being fine-tuned on SuperGLUE tasks. For example, MNLI from the GLUE benchmark (Wang et al., 2019) is used as an intermediate task for CB, RTE, and BoolQ (Wang et al., 2020). In contrast, our experiments with LLM did not use this method. Our models were only fine-tuned on the datasets in the SuperGLUE benchmark.
It is worth noting that on MultiRC, few-shot prompting performs slightly worse than zero-shot prompting. This may be because MultiRC involves long input contexts, and incorporating examples in a few-shot prompt can cause the total input length to approach or exceed the LLMs maximum context window.
Appendix G Evaluation on Reading Comprehension Tasks
We evaluate LLAMA2-7B-chat-hf on the SQuAD reading comprehension task, where the objective is to select a passage from a given context that best answers a question. Two fine-tuning strategies are compared: SFT, which directly uses the ground-truth answer as the training label, and PPO, which leverages reward functions based on Exact Match (EM) and F1 Score. EM metric is computed by comparing a normalized prediction against the normalized ground truth (with normalization involving lowercasing and punctuation removal); a perfect match yields an EM score of 1, otherwise 0. F1 score measures word-level overlap, balancing how many predicted words are correct (precision) and how many ground-truth words are included (recall).
Method EM F1
Non-fine-tuned 7.66 32.27
SFT 59.17 76.48
PPO 65.74 81.82
Dataset LLaMA2-7B-chat-hf LLaMA2-7B-base
CoLA 64.54 57.29
MRPC 88.41 80.35
STS-B 88.57 86.81
QQP 83.34 84.39
QNLI 93.17 92.95
MNLI-matched 88.63 87.15
MNLI-unmatched 88.95 88.22
RTE 84.90 83.75
SST-2 96.10 96.56
WNLI 74.65 67.61
Average 85.13 82.51
Table 12: Performance comparison of PPO fine-tuning on LLaMA2-7B-chat-hf and LLaMA2-7B-base across GLUE tasks. Best results for each dataset are highlighted in bold.
Appendix H Comparison of LLaMA2-7B-base and LLaMA2-7B-chat-hf on GLUE Tasks
We fine-tuned LLaMA2-7B-base with PPO and compared it to PPO-tuned LLaMA2-7B-chat-hf on GLUE tasks. The instruction-tuned chat model, with enhanced prompt-following capabilities, outperforms the base model on most tasks, averaging 85.13 vs. 82.51. This highlights the benefit of starting from an instruction-tuned checkpoint for task-specific RL.